OpenStudy (anonymous):

factor completely. r^2-22r+121

OpenStudy (anonymous):

how did you get that

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is actually \[r^2-22x+121=(r-11)^2\]

OpenStudy (misty1212):

you get it because \(121=11^2\) and \(-22=-2\times 11\)

OpenStudy (misty1212):

in general \[(a+b)^2=a^2+2ab+b^2\] so for example \[x^2-10x+25=(x-5)^2\]

OpenStudy (zkoqmosii):

You need to find out what two numbers add to equal -22 and multiply to equal 121. In this case it's -11 and -11. After finding the two numbers, you write it like this: (r-11)(r-11). You can also use the quadratic formula to find the solution.